Nicolas works in a lab. He needs a 15% acid solution for a certain test, but the supplier only ships a 10% solution and a 30% solution. He needs 10 liters of the 15% acid solution. How many liters of 10% solution and 30% solution should he use?
Answer by TimothyLamb(4379)

w = weak solution
s = strong solution
---
w + s = 10
s = 10 - w
10w + 30s = 15(10)
NOTE: the problem can also be solved with this equation (both equations will work):
0.10w + 0.30s = 0.15(10)
---
10w + 30s = 150
10w + 30(10 - w) = 150
10w + 300 - 30w = 150
20w = 150
---
w = 150/20
w = 7.5 liters
s = 2.5 liters
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php